# Wealth Thermalization Hypothesis

### v2

@@ Abstract
Figure S4: Comparison of the Lorenz curves for DE 2010 (black) and FR 2010 (blue dashed) with those of the RJS model (red curves; $N=10000$). The data of DE and FR were extracted from Ref. [32]. As in Fig. 3 the Gini coefficients $G=0.758$ (DE) and $G=0.679$ (FR) were used to determine the $\varepsilon$ values of the RJS model as $\varepsilon=0.1220$ (DE) and $\varepsilon=0.1659$ (FR) to match the Gini coefficients of the reference data.
@@
In this section, we present additional data. First Fig. S4 shows the Lorenz curves from Germany and France and the corresponding curves of RJS model (with matching Gini coefficients). These data were extracted from Ref. [32] with best possible precision and correspond to the period of 2010. The agreement with the RJS is comparable (not perfect but still rather close) as with the cases of US and World in Fig. 3. The Gini coefficients of Germany ($G=0.758$) and France ($G=0.679$) are both intermediate between UK ($G\approx 0.62$) and US/World ($G\approx 0.85/0.84$).
@@
Figure S5: As panel (b) of Fig. 4 but with a zoomed representation for $h\in[0.4,1]$ and $w\in[0,0.5]$ to increase the visibility between the close curves for US 2019 and World 2021.
@@
The next Fig. S5 shows a zoomed representation of panel (b) of Fig. 4 for $h\in[0.4,1]$ and $w\in[0,0.5]$ to increase the visibility between the close curves for US 2019 and World 2021 and to also to enhance the small differences with respect to the DL model (red lines).
[…]
@@
Furthermore, Fig. S6 presents results for the EQI model with $E_{0}>0$, $E_{m}=E_{0}+m/N$ and $\varepsilon=(E-E_{0})/(E_{N-1}-E_{0})\approx E-E_{0}$. In this case, the finite value $E_{0}>0$ induces an initial finite slope $E_{0}/E=E_{0}/(E_{0}+\varepsilon)$ in the Lorenz curve. We have verified that for the four cases shown in Fig. S6 this formula indeed represents the initial slope (see figure caption for the values). In this model, even the poorest households own a significant fraction of the wealth which is given by this slope. Here the range of possible Gini coefficients is quite limited with maximum values of $G_{\rm max}\approx 0.1$ or $0.4$ for $E_{0}=1$ or $E_{0}=0.1$ respectively. Due this reason it is not possible to match the data of US, UK, World etc. (with much larger Gini coefficients) to this model (for the cases shown in Figs. 6 and S6).
@@
Figure S7: Color plot of Lorenz curves for different models in the same style of Figure 2. The panels from top to bottom correspond to the DL model with parameter $a=16$ (1st row), to the shifted RMT semi-circle spectrum (2nd row), to the EQI model with given offset $E_{0}=0.1$ (3rd row) and $E_{0}=1$ (4th row). For the EQI model the rescaled energy $\varepsilon\in[0,0.5]$ for the vertical axis is given by $\varepsilon=(E-E_{0})/(E_{N-1}-E_{0})$ (same expression for the other models but with $E_{0}=0$). All cases correspond to $N=10000$ levels except for RMT with $N=1000$.
@@
Finally, Fig. S7 shows several color plots in the same style as Fig. 2, i.e. the color value (visible in the color bar) shows $w$ of the Lorenz curve as a function of $h$ ($x$-axis) and $\varepsilon$ ($y$-axis). The 2nd panel for the RMT model is rather similar to Fig. 2 for the RJS model, with a slight tendency for smaller $G$ values for $\varepsilon>0.2$ (at given $\varepsilon$, see also Fig. 5). The first panel for the DL model with $a=16$ ha a stronger condensation effect (i.e. more poor or poorer households) at $\varepsilon\approx 0.08$ as compared to $\varepsilon\approx 0.03$. Both bottom panels correspond to the EQI model with $E_{0}>0$ (here $\varepsilon=(E-E_{0})/(E_{N-1}-E_{0})$) with reduced Gini coefficients and where poor people own a significant fraction of the wealth, even at small values of $\varepsilon$.
III III. Analytical results for RJS model
@@
For the RJS model with finite $\varepsilon$ it is possible to obtain explicit formulas in the limit $N\to\infty$ by replacing the sums over $m$ with integrals over an energy variable $\tilde{E}=m/N\in[0,1]$. In the following, we also use $\varepsilon=E$ (since $E_{0}=0$ and $B=(N-1)/N\to 1$ for $N\to\infty$). In the limit $N\to\infty$ the implicit equation (S.3) becomes:
[…]
@@
$\mu=-(1-\mu)e^{-1/(\varepsilon-\mu)}\ .$ (S.5)
Both equations determine $\mu$ as a function of $\varepsilon\in]0,1[$. In the limit of small $\varepsilon$ one can simply iterate Eq. (S.5) by inserting $\mu_{0}=0$ in the RHS which gives $\mu_{1}=-e^{-1/\varepsilon}$ on the LHS which can be inserted in the RHS to obtain a better value $\mu_{2}$ etc. This procedure converges nicely for small $\varepsilon$ and for other values of $\varepsilon$ one can use standard techniques to solve these equations numerically and efficiently. For $\varepsilon\ll 1$, the first approximation $(-\mu)\approx e^{-1/\varepsilon}\ll\varepsilon$ is already very good.
To understand the limit of $|\mu|\gg 1$ it is more useful to consider $\varepsilon$ as a function of $\mu$ which is determined by (S.4). Expanding the logarithm in (S.4) up to 3rd order in $1/\mu$ one finds that
$\varepsilon\approx\frac{1}{2}\left(1+\frac{1}{6\mu}\right)\to\frac{1}{2}$ (S.6)
for $|\mu|\to\infty$ which is expected from the curve of $\mu$ in Fig. S1. The $1/\mu$ correction in (S.6) will be useful below.
@@
[…]
$=(\varepsilon-\mu)\ln\left(\frac{s-\mu}{-\mu}\right)$ (S.7)
$\quad\Rightarrow\quad s(h)$ $=(-\mu)\left(e^{h/(\varepsilon-\mu)}-1\right)$ (S.8)
@@
[…]
$=\frac{1}{\varepsilon}\bigg{[}(\varepsilon-\mu)s+\mu(\varepsilon-\mu)\ln\left(\frac{s-\mu}{-\mu}\right)\bigg{]}\ .$ (S.9)
Inserting (S.7) and (S.8) in (S.9) we obtain the following analytical expressions for the Lorenz curve:
$w(h)=\frac{-\mu}{\varepsilon}\left((\varepsilon-\mu)\left(e^{h/(\varepsilon-\mu)}-1\right)-h\right)$ (S.10)
$=\frac{1-\mu}{\varepsilon}\,e^{-1/(\varepsilon-\mu)}\left((\varepsilon-\mu)\left(e^{h/(\varepsilon-\mu)}-1\right)-h\right)\ .$ (S.11)
Here (S.11) has been obtained by replacing the global factor $\mu$ with (S.5) which gives a more convenient expression. Using (S.5), one can verify that (S.10) (and therefore also (S.11)) satisfy the conditions $w(0)=0$ and $w(1)=1$.
The expression (S.11) allows to take the limit $\varepsilon\ll 1$ with $\mu\approx-e^{-1/\varepsilon}\ll\varepsilon$ such that for $\varepsilon\ll 1$ we have the simplified Lorenz curve (replacing $\mu=0$ in (S.11)):
$w(h)$ $\approx e^{-1/\varepsilon}\left(e^{h/\varepsilon}-1-\frac{h}{\varepsilon}\right)\approx e^{(h-1)/\varepsilon}\ .$ (S.12)
@@
We have verified that both expressions (S.10) and (S.11) coincide with the numerical data shown in Fig. 1 up to graphical precision with an error below $10^{4}$ and for all values of $\varepsilon$ used in Fig. 1. The approximate formulas (S.12) are valid for $\varepsilon\leq 2$ with an error $\sim 10^{-2}$ for $\varepsilon=0.2$ (and smaller errors for smaller values of $\varepsilon$). This can be seen in Fig. S8 comparing the data for $\varepsilon=0.1,0.2,0.3$ between the analytic expressions and the data for $N=10000$. Even for $\varepsilon=0.3$ only a modest deviation of the approximate curve is visible while here and in all other cases the more precise expression (S.11) matches the numerical data very closely.
@@
$w(h)$ $=\frac{(-\mu)}{2\varepsilon(\varepsilon-\mu)}\,h^{2}\ .$ (S.13)
We know that the limit $|\mu|\to\infty$ corresponds to $\varepsilon\to 1/2$ and in this case (S.13) is valid for all $h\in[0,1]$. This gives the very simple formula $w=h^{2}$ (which is also obvious from the fact that $\rho_{m}=1/N=$ const. for $|\mu|\to\infty$ and the way the Lorenz curve is constructed from $\rho_{m}$).
Figure S8: Comparison of Lorenz curves of wealth fraction $w$ versus household fraction $h$ for the analytical model with the numerical data of the RJS model for finite $N=10000$ and for three key values of the rescaled energy $\varepsilon=0.1,0.2,0.3$ (top to bottom). Left panels shows the difference between the analytical model and numerical data and right panels show directly the curves $w$ versus $h$ for the numerical data (red lines and plus symbols) and the analytical model. Blue lines/data points correspond to the formula (S.11) valid for all values of $\varepsilon$ and using the appropriate value of the chemical potential $\mu$ determined by the implicit equation (S.4). Green lines/data points correspond to the (second) approximate formula (S.12) valid for small $\varepsilon\leq 0.2$. The discrete points of data in the top right panel for $\varepsilon=0.1$ at values close to $w=0$ indicate finite values for $\rho_{0}=0.1129$, $\rho_{0}+\rho_{1}=0.1660$, etc. which are due to RJ condensation.
@@
[…]
$=1+\frac{2\mu}{\varepsilon}\left[(\varepsilon-\mu)^{2}(e^{1/(\varepsilon-\mu)}-1)-(\varepsilon-\mu)-\frac{1}{2}\right]$ (S.14)
$=1-\frac{\mu}{\varepsilon}-2(\varepsilon-\mu)\ .$ (S.15)
@@
Here the second simpler expression (S.15) has been obtained by replacing the exponential term in (S.14) using the implicit equation of $\mu$. The limit $\varepsilon\ll 1$ with $\mu\approx-e^{-1/\varepsilon}$ gives $G\approx 1-2\varepsilon$ which matches well the values of $G$ given in the caption of Fig. 1 for $\varepsilon\leq 0.1$ (rather close value for $\varepsilon=0.2$). The other values are matched exactly by the more precise expression (S.15). Furthermore, inserting the expression (S.6) for large $|\mu|$ in (S.15) one finds (confirms) that $G=1/3$ for $\varepsilon=1/2$ (here it is necessary to keep the $1/\mu$ correction in (S.6) to obtain the correct result for $G$).
@@
Using the analytical expression (S.11) for $w(h)$ it is also straightforward to compute (with simple numerics) the inverse function $h(w)$. Using this and the analytical expression (S.15) for the Gini coefficient, we have also recomputed the curves for $G(\varepsilon)$, $h(2\%)$, $1-h(25\%)$ (both as a function of $\varepsilon$) and verified that the analytical curves coincide with the numerical curves shown in Figs. 5 and 6 (for the RJS model at $N=10000$) up to graphical precision (typical error $\sim 10^{-4}$).
@@
One might be concerned that the integral approximation is not very good for small $\mu$ (close to the singularity of the first term in (S.3)) and some finite but large value of $N$ such as $N=10000$. This is true but the integral provides a modified logarithmic singularity which allows also to mimic correctly the condensation effect with correct probabilities. Therefore even though the values of $\mu$ are modified for $\varepsilon\ll 1$ (but still $0<-\mu\ll\varepsilon\ll 1$ for both models !) the resulting probabilities (e.g. integrals or sums of $\rho_{m}$ over some interval in $\tilde{E}=m/N$) are the same. The values of $\mu$ obtained by the continuous analytical model match very well the curve shown in Fig. S1 but of course this figure does not allow to verify if $\mu\approx-e^{-1/\varepsilon}$ (continuous model) or $\mu\approx-\varepsilon/(N-1)$ (for the finite $N$ model with discrete sums) which are both close to zero in the figure. In any case, we find that the analytical expressions given here (if $\mu$ is properly evaluated by its implicit equation (S.4) and if properly evaluated by avoiding numerical instabilities of some formulas in some special cases) match the numerical data with an error that scales with $1/N$.
@@
Without going into details, we mention that we have also considered a more refined version of the continuous model using a finite value of $N$ and keeping the first singular term separate from the integral (which starts at $s=1/N$ and not $S=0$). In this case, we obtain a modified implicit equation of $\mu$ which results in values of $\mu$ closer to the model of finite $N$ but the resulting physical quantities ($w(h)$ curves, Gini coefficients etc.) are (numerically with an error $<10^{-4}$) the same as both the numerical data and the simple model. The resulting analytical expressions of the refined model are slightly modified (essentially replacing $h$ by $h-\rho_{0}$ for $h\geq\rho_{0}$ in the formula of the Lorenz curve and using $w(h)=0$ for $h<\rho_{0}$ where $\rho_{0}$ may now have a finite value). Note that the initial interval $h\in[0,\rho_{0}[$ with exactly $w(h)=0$ for the refined and also the discrete model translates to exponentially small values $w(h)\approx h^{2}\,e^{-1/\varepsilon}/(2\varepsilon)$ for the simple analytical model (replacing $\mu\approx-e^{-1/\varepsilon}$ in (S.13)).
@@
IV IV. Data for companies of stock exchange at New York, London, Hong Kong

### v3

@@ A.2 Additional data
Figure A.4: Comparison of the Lorenz curves for DE 2010 (black) and FR 2010 (blue dashed) with those of the RJS model (red curves; $N=10000$). The data of DE and FR were extracted from Ref. [56]. As in Fig. I.3 the Gini coefficients $G=0.758$ (DE) and $G=0.679$ (FR) were used to determine the $\varepsilon$ values of the RJS model as $\varepsilon=0.1220$ (DE) and $\varepsilon=0.1659$ (FR) to match the Gini coefficients of the reference data.
@@
In this section, we present additional data. First Fig. A.4 shows the Lorenz curves from Germany and France and the corresponding curves of RJS model (with matching Gini coefficients). These data were extracted from Ref. [56] with best possible precision and correspond to the period of 2010. The agreement with the RJS is comparable (not perfect but still rather close) as with the cases of US and World in Fig. I.3. The Gini coefficients of Germany ($G=0.758$) and France ($G=0.679$) are both intermediate between UK ($G\approx 0.62$) and US/World ($G\approx 0.85/0.84$).
@@
Figure A.5: As panel (b) of Fig. I.4 but with a zoomed representation for $h\in[0.4,1]$ and $w\in[0,0.5]$ to increase the visibility between the close curves for US 2019 and World 2021.
@@
The next Figure A.5 shows a zoomed representation of panel (b) of Figure I.4 for $h\in[0.4,1]$ and $w\in[0,0.5]$ to increase the visibility between the close curves for US 2019 and World 2021 and to also to enhance the small differences with respect to the DL model (red lines).
Figure A.6: Lorenz curves of the thermalized EQI model ($N=10000$) with the two offset values $E_{0}=0.1$ and $E_{0}=1$ and for each case for two values of the rescaled energy $\varepsilon=(E-E_{0})/(E_{N-1}-E_{0})$. Note that for $E_{0}=1\Rightarrow E_{N-1}\approx 2E_{0}$ and for $E_{0}=0.1\Rightarrow E_{N-1}\approx 11E_{0}$. The dashed line corresponds to the line of perfect equipartition $w=h$. The Gini coefficients $G$ for all curves are G=0.4239, 0.3000, 0.1162, 0.04286G=0.4239,\,0.3000,\,0.1162,\,0.04286 (bottom to top). These curves show a finite initial slope with value $E_{0}/(\varepsilon+E_{0})=0.4,\,0.6667,\,0.7143,\,0.9524$ (bottom to top).
@@
Furthermore, Fig. A.6 presents results for the EQI model with $E_{0}>0$, $E_{m}=E_{0}+m/N$ and $\varepsilon=(E-E_{0})/(E_{N-1}-E_{0})\approx E-E_{0}$. In this case, the finite value $E_{0}>0$ induces an initial finite slope $E_{0}/E=E_{0}/(E_{0}+\varepsilon)$ in the Lorenz curve. We have verified that for the four cases shown in Fig. A.6 this formula indeed represents the initial slope (see figure caption for the values). In this model, even the poorest households own a significant fraction of the wealth which is given by this slope. Here the range of possible Gini coefficients is quite limited with maximum values of $G_{\rm max}\approx 0.1$ or $0.4$ for $E_{0}=1$ or $E_{0}=0.1$ respectively. Due this reason it is not possible to match the data of US, UK, World etc. (with much larger Gini coefficients) to this model (for the cases shown in Figs. I.6A.6).
@@
Figure A.7: Color plot of Lorenz curves for different models in the same style of Figure I.2. The panels correspond to the DL model with parameter $a=16$ (top left), to the shifted RMT semi-circle spectrum (top right), to the EQI model with given offset $E_{0}=0.1$ (bottom left) and $E_{0}=1$ (bottom right). For the EQI model the rescaled energy $\varepsilon\in[0,0.5]$ for the vertical axis is given by $\varepsilon=(E-E_{0})/(E_{N-1}-E_{0})$ (same expression for the other models but with $E_{0}=0$). All cases correspond to $N=10000$ levels except for RMT with $N=1000$.
@@
Finally, Fig. A.7 shows several color plots in the same style as Fig. I.2, i.e. the color value (visible in the color bar) shows $w$ of the Lorenz curve as a function of $h$ ($x$-axis) and $\varepsilon$ ($y$-axis). The 2nd panel for the RMT model is rather similar to Fig. I.2 for the RJS model, with a slight tendency for smaller $G$ values for $\varepsilon>0.2$ (at given $\varepsilon$, see also Fig. I.5). The first panel for the DL model with $a=16$ has a stronger condensation effect (i.e. more poor or poorer households) at $\varepsilon\approx 0.08$ as compared to $\varepsilon\approx 0.03$. Both bottom panels correspond to the EQI model with $E_{0}>0$ (here $\varepsilon=(E-E_{0})/(E_{N-1}-E_{0})$) with reduced Gini coefficients and where poor people own a significant fraction of the wealth, even at small values of $\varepsilon$.
A.3 Analytical results for RJS model
@@
For the RJS model with finite $\varepsilon$ it is possible to obtain explicit formulas in the limit $N\to\infty$ by replacing the sums over $m$ with integrals over an energy variable $\tilde{E}=m/N\in[0,1]$. In the following, we also use $\varepsilon=E$ (since $E_{0}=0$ and $B=(N-1)/N\to 1$ for $N\to\infty$). In the limit $N\to\infty$ the implicit equation (A.3) becomes:
$1$ $=(\varepsilon-\mu)\int_{0}^{1}\frac{1}{\tilde{E}-\mu}\,d\tilde{E}=(\varepsilon-\mu)\ln\left(\frac{1-\mu}{-\mu}\right)$ (A.4)
@@
$\mu=-(1-\mu)e^{-1/(\varepsilon-\mu)}\ .$ (A.5)
Both equations determine $\mu$ as a function of $\varepsilon\in]0,1[$. In the limit of small $\varepsilon$ one can simply iterate Eq. (A.5) by inserting $\mu_{0}=0$ in the RHS which gives $\mu_{1}=-e^{-1/\varepsilon}$ on the LHS which can be inserted in the RHS to obtain a better value $\mu_{2}$ etc. This procedure converges nicely for small $\varepsilon$ and for other values of $\varepsilon$ one can use standard techniques to solve these equations numerically and efficiently. For $\varepsilon\ll 1$, the first approximation $(-\mu)\approx e^{-1/\varepsilon}\ll\varepsilon$ is already very good.
To understand the limit of $|\mu|\gg 1$ it is more useful to consider $\varepsilon$ as a function of $\mu$ which is determined by (A.4). Expanding the logarithm in (A.4) up to 3rd order in $1/\mu$ one finds that
$\varepsilon\approx\frac{1}{2}\left(1+\frac{1}{6\mu}\right)\to\frac{1}{2}$ (A.6)
for $|\mu|\to\infty$ which is expected from the curve of $\mu$ in Fig. A.1. The $1/\mu$ correction in (A.6) will be useful below.
@@
$\phantom{\quad\Rightarrow\quad}h(s)$ $=(\varepsilon-\mu)\int_{0}^{s}\frac{1}{\tilde{E}-\mu}\,d\tilde{E}=(\varepsilon-\mu)\ln\left(\frac{s-\mu}{-\mu}\right)$ (A.7)
$\quad\Rightarrow\quad s(h)$ $=(-\mu)\left(e^{h/(\varepsilon-\mu)}-1\right)$ (A.8)
@@
$w(s)$ $=\frac{\varepsilon-\mu}{\varepsilon}\int_{0}^{s}\frac{\tilde{E}}{\tilde{E}-\mu}\,d\tilde{E}$
$=\frac{1}{\varepsilon}\bigg[(\varepsilon-\mu)s+\mu(\varepsilon-\mu)\ln\left(\frac{s-\mu}{-\mu}\right)\bigg]\ .$ (A.9)
Inserting (A.7) and (A.8) in (A.9) we obtain the following analytical expressions for the Lorenz curve:
$w(h)$ $=\frac{-\mu}{\varepsilon}\left((\varepsilon-\mu)\left(e^{h/(\varepsilon-\mu)}-1\right)-h\right)$ (A.10)
$=\frac{1-\mu}{\varepsilon}\,e^{-1/(\varepsilon-\mu)}\left((\varepsilon-\mu)\left(e^{h/(\varepsilon-\mu)}-1\right)-h\right)\ .$ (A.11)
Here (A.11) has been obtained by replacing the global factor $\mu$ with (A.5) which gives a more convenient expression. Using (A.5), one can verify that (A.10) (and therefore also (A.11)) satisfy the conditions $w(0)=0$ and $w(1)=1$.
The expression (A.11) allows to take the limit $\varepsilon\ll 1$ with $\mu\approx-e^{-1/\varepsilon}\ll\varepsilon$ such that for $\varepsilon\ll 1$ we have the simplified Lorenz curve (replacing $\mu=0$ in (A.11)):
$w(h)$ $\approx e^{-1/\varepsilon}\left(e^{h/\varepsilon}-1-\frac{h}{\varepsilon}\right)\approx e^{(h-1)/\varepsilon}\ .$ (A.12)
@@
We have verified that both expressions (A.10) and (A.11) coincide with the numerical data shown in Fig. I.1 up to graphical precision with an error below $10^{-4}$ and for all values of $\varepsilon$ used in Fig. I.1. The approximate formulas (A.12) are valid for $\varepsilon\leq 2$ with an error $\sim 10^{-2}$ for $\varepsilon=0.2$ (and smaller errors for smaller values of $\varepsilon$). This can be seen in Fig. A.8 comparing the data for $\varepsilon=0.1,0.2,0.3$ between the analytic expressions and the data for $N=10000$. Even for $\varepsilon=0.3$ only a modest deviation of the approximate curve is visible while here and in all other cases the more precise expression (A.11) matches the numerical data very closely.
@@
$w(h)$ $=\frac{(-\mu)}{2\varepsilon(\varepsilon-\mu)}\,h^{2}\ .$ (A.13)
We know that the limit $|\mu|\to\infty$ corresponds to $\varepsilon\to 1/2$ and in this case (A.13) is valid for all $h\in[0,1]$. This gives the very simple formula $w=h^{2}$ (which is also obvious from the fact that $\rho_{m}=1/N=$ const. for $|\mu|\to\infty$ and the way the Lorenz curve is constructed from $\rho_{m}$).
Figure A.8: Comparison of Lorenz curves of wealth fraction $w$ versus household fraction $h$ for the analytical model with the numerical data of the RJS model for finite $N=10000$ and for three key values of the rescaled energy $\varepsilon=0.1,0.2,0.3$ (top to bottom). Left panels shows the difference between the analytical model and numerical data and right panels show directly the curves $w$ versus $h$ for the numerical data (red lines and plus symbols) and the analytical model. Blue lines/data points correspond to the formula (A.11) valid for all values of $\varepsilon$ and using the appropriate value of the chemical potential $\mu$ determined by the implicit equation (A.4). Green lines/data points correspond to the (second) approximate formula (A.12) valid for small $\varepsilon\leq 0.2$. The discrete points of data in the top right panel for $\varepsilon=0.1$ at values close to $w=0$ indicate finite values for $\rho_{0}=0.1129$, $\rho_{0}+\rho_{1}=0.1660$, etc. which are due to RJ condensation.
@@
$G$ $=1-2\int_{0}^{1}w(h)\,dh$
$=1+\frac{2\mu}{\varepsilon}\left[(\varepsilon-\mu)^{2}(e^{1/(\varepsilon-\mu)}-1)-(\varepsilon-\mu)-\frac{1}{2}\right]$ (A.14)
$=1-\frac{\mu}{\varepsilon}-2(\varepsilon-\mu)\ .$ (A.15)
@@
Here the second simpler expression (A.15) has been obtained by replacing the exponential term in (A.14) using the implicit equation of $\mu$. The limit $\varepsilon\ll 1$ with $\mu\approx-e^{-1/\varepsilon}$ gives $G\approx 1-2\varepsilon$ which matches well the values of $G$ given in the caption of Fig. I.1 for $\varepsilon\leq 0.1$ (rather close value for $\varepsilon=0.2$). The other values are matched exactly by the more precise expression (A.15). Furthermore, inserting the expression (A.6) for large $|\mu|$ in (A.15) one finds (confirms) that $G=1/3$ for $\varepsilon=1/2$ (here it is necessary to keep the $1/\mu$ correction in (A.6) to obtain the correct result for $G$).
@@
Using the analytical expression (A.11) for $w(h)$, it is also straightforward to compute (with simple numerics) the inverse function $h(w)$. Using this and the analytical expression (A.15) for the Gini coefficient, we have also recomputed the curves for $G(\varepsilon)$, $h(2\%)$, $1-h(25\%)$ (both as a function of $\varepsilon$) and verified that the analytical curves coincide with the numerical curves shown in Fig. I.5 and Fig. I.6 (for the RJS model at $N=10000$) up to graphical precision (typical error $\sim 10^{-4}$).
@@
One might be concerned that the integral approximation is not very good for small $\mu$ (close to the singularity of the first term in (A.3)) and some finite but large value of $N$ such as $N=10000$. This is true but the integral provides a modified logarithmic singularity which allows also to mimic correctly the condensation effect with correct probabilities. Therefore even though the values of $\mu$ are modified for $\varepsilon\ll 1$ (but still $0<-\mu\ll\varepsilon\ll 1$ for both models !) the resulting probabilities (e.g. integrals or sums of $\rho_{m}$ over some interval in $\tilde{E}=m/N$) are the same. The values of $\mu$ obtained by the continuous analytical model match very well the curve shown in Fig. A.1 but of course this figure does not allow to verify if $\mu\approx-e^{-1/\varepsilon}$ (continuous model) or $\mu\approx-\varepsilon/(N-1)$ (for the finite $N$ model with discrete sums) which are both close to zero in the figure. In any case, we find that the analytical expressions given here (if $\mu$ is properly evaluated by its implicit equation (A.4) and if properly evaluated by avoiding numerical instabilities of some formulas in some special cases) match the numerical data with an error that scales with $1/N$.
@@
Without going into details, we mention that we have also considered a more refined version of the continuous model using a finite value of $N$ and keeping the first singular term separate from the integral (which starts at $s=1/N$ and not $S=0$). In this case, we obtain a modified implicit equation of $\mu$ which results in values of $\mu$ closer to the model of finite $N$ but the resulting physical quantities ($w(h)$ curves, Gini coefficients etc.) are (numerically with an error $10^{-4}$) the same as both the numerical data and the simple model. The resulting analytical expressions of the refined model are slightly modified (essentially replacing $h$ by $h-\rho_{0}$ for $h\geq\rho_{0}$ in the formula of the Lorenz curve and using $w(h)=0$ for $h<\rho_{0}$ where $\rho_{0}$ may now have a finite value). Note that the initial interval $h\in[0,\rho_{0}[$ with exactly $w(h)=0$ for the refined and also the discrete model translates to exponentially small values $w(h)\approx h^{2}\,e^{-1/\varepsilon}/(2\varepsilon)$ for the simple analytical model (replacing $\mu\approx-e^{-1/\varepsilon}$ in (A.13)).
@@
A.4 Data for companies of stock exchange at New York, London, Hong Kong